OpenStudy (anonymous):

one number is 5 more than twice another. the sum is -22

OpenStudy (anonymous):

x = 2y +5 x + y = -22 2y + 5 + y = -22 x = 2(-9) + 5 3y = -27 x= -13 y = -9